Grace;s average driving speed is 3 mph faster than Thelma's. In the same length of time it takes Grace to drive 156 miles, Thelma only drives 147 miles. What is Grace's average speed?
Answer by checkley71(8403)

156/(X+3)=147/X
156X=147X+441
156X-147X=441
9X=441
X=441/9
X=45 MPH
X+3=45+3=48 MPH FOR GRACE